Documenta Mathematica

Band 9, 2004

Mike Develin and Bernd Sturmfels

Tropical Convexity  1–27

Serge Yagunov

Rigidity II: Non-Orientable Case  29–40

V. Franjou and T. Pirashvili

Comparison of Abelian Categories Recollements  41–56

Rupert L. Frank and Roman G. Shterenberg On the Scattering Theory of the Laplacian with a Periodic Boundary Condition.

II. Additional Channels of Scattering  57–77

Paul S. Muhly and Mark Tomforde

Adding Tails to C^{∗}-Correspondences  79–106

N. Filonov and F. Klopp Absolute Continuity of the Spectrum of a Schrödinger Operator with a Potential Which is Periodic

in Some Directions and Decays in Others  107–121

Gabor Wiese Dihedral Galois Representations

and Katz Modular Forms  123–133

N. Filonov and F. Klopp Erratum to the paper “Absolute Continuity of the Spectrum of a Schrödinger Operator with a Potential Which is Periodic

in Some Directions and Decays in Others”  135–136

William Arveson

The Free Cover of a Row Contraction  137–161

A. Polishchuk Classification of Holomorphic Vector Bundles

on Noncommutative Two-Tori  163–181

R. Preeti and J.-P. Tignol

Multipliers of Improper Similitudes  183–204

Mike Develin and Bernd Sturmfels

Erratum for “Tropical Convexity”  205–206

Xiaochun Liu and Ingo Witt Asymptotic Expansions for Bounded Solutions

to Semilinear Fuchsian Equations  207–250

Dennis Snow Bounds for the Anticanonical Bundle of a

Homogeneous Projective Rational Manifold  251–263

E. Ballico, C. Casagrande, and C. Fontanari

Moduli of Prym Curves  265–281

A. Ancona, B. Helffer, T. Hoffmann-Ostenhof

Nodal Domain Theorems à la Courant  283–299

Louis Mahé, Ján Mináč, and Tara L. Smith Additive Structure of Multiplicative Subgroups

of Fields and Galois Theory  301–355

David Burns On the Values of Equivariant Zeta Functions

of Curves over Finite Fields  357–399

Wolfgang Kühnel

Tight Embeddings of Simply Connected 4-Manifolds  401–412

Adriano Marmora

Irrégularité et Conducteur de Swan p-adiques  413–433

Yasuhito Miyamoto On Connecting Orbits of

Semilinear Parabolic Equations on S^{1}  435–469

Ekaterina Amerik Some Remarks on Morphisms

between Fano Threefolds  471–486

Alexander Nenashev Projective Bundle Theorem

in Homology Theories with Chern Structure  487–497

Pierre-Emmanuel Chaput Theoremes d’Annulation et Lieux

de Degenerescence en Petit Corang  499–525

Colette Moeglin Stabilité en Niveau 0,

pour les Groupes Orthogonaux Impairs p-Adiques  527–564

Jean-Louis Tu Non-Hausdorff Groupoids,

Proper Actions and K-Theory  565–597

Konstantin Ardakov The Centre of Completed Group Algebras

of Pro-p Groups  599–606

Liu Lanzhe Some Sharp Weighted Estimates

for Multilinear Operators  607–622

Nicolas Perrin Rational Curves on

Homogeneous Cones  623–637

Toke Meier Carlsen and Søren Eilers Matsumoto K-Groups

Associated to Certain Shift Spaces  639–671

vi